OpenStudy (anonymous):

PLEASE HELP ME~ If Earth's mass and radius both suddenly doubled, what would be the new value of the acceleration of gravity near Earth's surface? (a) 9.80 m/s2 (b) 4.90 m/s2 (c) 2.45 m/s2 (d) 19.6 m/s2 (e) 12.6 m/s2

OpenStudy (anonymous):

\[a=\frac{ Gm }{ r^2 }\]Double mass and radius to get\[a' = \frac{ G(2m) }{ \left( 2r \right)^2 } = \frac{2Gm}{4r^2}=\frac{Gm}{2r^2}\]Compare the two.

OpenStudy (anonymous):

@ospreytriple what do you think the answer? Cos I don't know too haha

OpenStudy (anonymous):

I won't tell you the answer, but compare the two equations above. what is the relationship between a and a' ?

OpenStudy (anonymous):

@ospreytriple I don't know :3